OpenStudy (anonymous):

Find any stationary points in the interval 0

OpenStudy (anonymous):

Please also explain to me how how you use the chain rule with this.

OpenStudy (aum):

\( \Large y = (\sin(x))^2 + 2\cos(x) \\ \Large \frac {dy}{dx} = \frac{d}{dx}(\sin(x))^2 + \frac {d}{dx}(2*\cos(x)) \\ \Large ~~~~ = 2(\sin(x))^1 * \frac {d}{dx}(\sin(x)) - 2\frac {d}{dx}\cos(x) \\ \Large ~~~~ = 2\sin(x)\cos(x) - 2\sin(x) \\ \Large ~~~~ = 2\sin(x)(\cos(x) - 1) \\ \) To find stationary points, set \(\Large \frac {dy}{dx} = 0 \) and solve for x.

OpenStudy (anonymous):

So one of the sin x's is treated differently from the other?

OpenStudy (aum):

Not sure of the question. But to differentiate (sin(x))^2 we first apply the power rule which says the derivative of (something)^n = n*(something^(n-1) and then we apply the chain rule by multiplying the above by derivative of (something). So d/dx (sin x)^2 = 2(sin x)^1 * d/dx(sin x) = 2sin(x)cos(x)

OpenStudy (anonymous):

I understand, thank you very much!

OpenStudy (aum):

If it would help you can make use of substitution: If y = (sin x)^2 find dy/dx. Let u = sin x y = u^2 dy/dx = dy/du * du/dx = d/du(u^2) * d/dx(sin x) = 2u^1 * cos(x) = 2sin(x)cos(x) (by putting u back as sin x) You are welcome.

OpenStudy (anonymous):

That's the first someone tried teaching me the substitution thing that I actually understood! Wow thanks and thanks again!

OpenStudy (aum):

Glad to be able to help. And you are very welcome. :)